OpenStudy (anonymous):

-13X=97-19Y -17X=83+19y

OpenStudy (anonymous):

\[=-30x=180\]\[x=-6\]\[-17(-6)=83+19y\]\[y=1\] Check \[=-13(-6)=97-19(1)\]\[78=78\]checks